Toxic pollutants. The annual cost in dollars for removing
p% of the toxic chemicals from a town’s water supply is
given by the formula
C(p)=500,000/100-p
a) Use the accompanying graph to estimate the cost for
removing 90% and 95% of the toxic chemicals.
b) Use the formula to find C(99.5) and C(99.9).
c) What happens to the cost as the percentage of
pollutants removed approaches 100%?

Answer by checkley77(12844)

C(p)=500,000/100-p
c(p)=500,000/(100-90)
c(p)=500,000/10
c(p)=50,000 ans.
--------------------
c(p)=500,000/(100-95)
c(p)=500,000/5
c(p)=100,000 ans.
----------------------
c(p)=500,000/(100-99.5)
c(p)=500,000/.5
c(p)=1,000,000 ans.
-----------------------
c(p)=500,000/(100-99.9)
c(p)=500,000/.1
c(p)=5,000,000 ans.
-----------------------
As the percentace approaches 100% the cost approaches infinity.